In order to conduct a hypothesis test for the population proportion, you sample 400 observations that result in 212 successes. (You may find it useful to reference the appropriate table: z table or t table)

H0: p ≥ 0.54; HA: p < 0.54.

a-1. Calculate the value of the test statistic. (Negative value should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)

H0: p = 0.54; HA: p ≠ 0.54.

b-1. Calculate the value of the test statistic. (Negative value should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)
Answer #1

Solution:

a-1 The value of the test statistic is:

P-p 2= p(1-P)

Where:

22 =0.53 400

  0.53 – 0.54 ..2= 0.54(1-0.54) 400

  - -0.40

b-1. Calculate the value of the test statistic.

Answer: The value of the test statistic is:

P-p 2= p(1-P)

Where:

22 =0.53 400

  0.53 – 0.54 ..2= 0.54(1-0.54) 400

  - -0.40